alright I have this problem and need step by step how to solve cause I am confused
the width of a rectangle is 4ftless then the length. The area is 10 ft squared. Find the length and the width.
The answer in the back of the book is
2 + radican 14 ft is length
radicand 14 - 2ft
I am not sure how to get this

Answer by scott8148(6628)

W + 4 = L

A = L * W

10 = (W + 4) * W = W^2 + 4W

0 = W^2 + 4W - 10

use quadratic formula to find W
___ then substitute into "L = W + 4" to find L